OpenStudy (anonymous):

Describe the differences between the graph of y = -3(x + 7)2 - 10 and the standard position graph of y = x2.

OpenStudy (anonymous):

y = mx + b y = (-3)((x + 7)^2) - (10)

OpenStudy (anonymous):

Do you know how to graph this?

OpenStudy (anonymous):

yes u find the vertex by this formula x=-b/2a

OpenStudy (anonymous):

then u take 2 points away from it

OpenStudy (anonymous):

I find it easier to use a function to find two separate points, want me to show you how?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y = -3(x + 7)^2 - 10 y = f(x) f(x) = -3(x + 7)^2 - 10 Let x = 0 f(0) = -3(0 + 7)^2 - 10 f(0) = -3(7)^2 - 10 f(0) = -3(49) - 10 f(0) = -147 - 10 f(0) = -157 So when x = 0 y = -157 (0,-157) Also, when x = 0, you have your y intercept when you find y. So the y intercept is -157.

OpenStudy (anonymous):

All you have to do is take the equation an plug in any value for x and solve for y. Then you have an ordered pair. You need two ordered pairs to graph it.

OpenStudy (anonymous):

ok